# GraphAU-Pain: Graph-based Action Unit Representation for Pain Intensity Estimation

### v1

@@ I INTRODUCTION
Recent advances in graph neural networks (GNNs), such as Multi-dimensional Edge Feature-based AU Relation Graph for AU (ME-GraphAU) [20] and Graph Relation Network (GRU) [25], have demonstrated promising AU prediction performance on the DISFA dataset [21] and BP4D dataset [31]. These methods use CNN backbones like ResNet [11] or VGG [14] to learn image features for each AU, then construct a relational graph based on the AU features. Each node in such graphs represents an AU, and each edge represents the relationship between a pair of AUs [17]. Through message propagation in GNN layers, the output AU features can capture individual AU information from neighbors as well as structural information [15]. Such graph-based AU features can then be aggregated to construct a full-face representation informed by AUs and their relationships for downstream tasks [16], inspiring us to incorporate graph-based AU detection into pain intensity estimation.
Motivated by the above, we introduce GraphAU-Pain and summarize our three contributions as follows:
  • GraphAU-Pain Model. We fuse a ResNet-50 backbone with graph-based AU representation learning, addressing the shortcomings of purely image-driven methods.
  • Cross-Dataset Transfer with Relabeling. To mitigate poor direct training on UNBC, we fine-tune GraphAU-Pain on DISFA-pretrained weights (sharing most AUs), partially relabeling UNBC to create UNBC+ to handle missing AU annotations and boost performance.
@@
II MATERIALS AND METHODS
II-A Implementation Overview
@@
The GraphAU-Pain model is designed to estimate pain intensity based on facial image data, and its training involves two key steps. First, the full-face and AU representation learning modules are trained for AU occurrence prediction. These modules are adapted from the AU Relationship-aware Node Feature Learning (ANFL) component of ME-GraphAU [20]. With only a $20\%$ average F1-score achieved when directly training ANFL for AU occurrence prediction on the UNBC dataset, we developed a cross-dataset transfer learning strategy to improve performance. This strategy transfers AU prediction capabilities from the DISFA dataset model trained by Luo et al. [20] to the UNBC dataset. To address differences in AU labels between datasets, we created a relabeled UNBC+ dataset to align the AU annotations and used undersampling to address AU imbalance. Second, with the weights trained for AU prediction, GraphAU-Pain is then trained for pain intensity estimation with the full UNBC dataset. Facial images are first processed through the CNN backbone to extract pixel-wise full-face representations. These representations are then transformed into graph-based structures to learn AU-specific features. Finally, the full-face and AU-based features are combined to estimate pain intensity. The performance of the model is evaluated using two metrics: average F1-score and accuracy.
@@
Figure 1: Illustration of the GraphAU-Pain model architecture.
A global CNN backbone (orange) extracts a full-face representation, which is split into three feature branches by the AU representation learning module (green) and processed by the pain level classification module (blue). ANFL denotes the AU Relationship-aware Node Feature Learning component, FC the fully-connected layer, and MatMult matrix multiplication.
II-B Fine-tuning Representation Learning Modules
Our preliminary experimental results showed that directly training ANFL on the UNBC dataset failed to provide accurate AU occurrence prediction. Similarly, directly training GraphAU-Pain on UNBC yielded unsatisfactory results, with $30\sim 40\%$ F1-score and $60\sim 80\%$ accuracy after trying several settings. To tackle this problem and improve AU prediction performance through transfer learning, in this work we fine-tuned ANFL’s weights pretrained on DISFA, which shares 6 out of 8 AUs with UNBC, 3 of them used in PSPI calculation. To relabel UNBC so its AU labels align with DISFA, the facial images of UNBC are passed to the pretrained ANFL module to obtain 8 predicted AU labels. Of these 8 labels, the 6 that exist in UNBC are replaced with the original values, while for AU1 and AU2 predicted values are used. We refer to this relabelled dataset as UNBC+ to distinguish it from the original UNBC dataset.
II-C Model
@@
After acquiring the backbone feature $\mathbf{h_{b}}$, the AU representation module is applied to learn two distinct embeddings: $\mathbf{h_{a}}\in\mathbb{R}^{d_{\text{AU}}}$, where $d_{\text{AU}}=8$, encoding each AU’s occurrence; and $\mathbf{h_{g}}\in\mathbb{R}^{d_{\text{node}}}$, where $d_{\text{node}}=512$, encoding the complete structure of the AU relational graph. The AU representation module first learns the AU relational graph $\mathcal{G}$ from $\mathbf{h_{b}}$ as in the original ANFL. It begins with applying $d_{\text{AU}}$ fully connected layers on $\mathbf{h_{b}}$ to obtain a set of initial AU representations $\mathbf{V}\in\mathbb{R}^{d_{\text{AU}}\times d_{\text{node}}}$ as node features. Each node then connects with its $K=3$ most similar nodes based on dot-product similarity. Next, messages are passed through a graph convolutional layer:
$\mathbf{V}^{\prime}\;=\;\text{ReLU}\Bigl{[}\mathbf{V}+\text{BN}\bigl{(}\mathbf{A}^{\top}\text{FC}_{1}(\mathbf{V})+\text{FC}_{2}(\mathbf{V})\bigr{)}\Bigr{]}\;\in\;\mathbb{R}^{d_{\text{AU}}\times d_{\text{node}}},$ (1)
where BN refers to batch normalization, FC refers to fully connected layers, and $A$ is the normalized adjacency matrix. Then, AU features $\mathbf{h_{a}}\in\mathbb{R}^{d_{\text{AU}}}$ are obtained by a similarity-calculating strategy [20], and the graph representation $\mathbf{h_{g}}\in\mathbb{R}^{d_{\text{node}}}$ is computed by a global sum pooling of the node embeddings. Although it would be possible to add an edge feature learning module here [20], we omitted it to avoid overfitting on the UNBC dataset. After obtaining these three features, three fully connected layers with ReLU activation map each of ${\mathbf{h_{b}}}$, ${\mathbf{h_{a}}}$, and ${\mathbf{h_{g}}}$ to a common dimension of $36$, producing ${\mathbf{h_{b}^{\prime}}}$, ${\mathbf{h_{a}^{\prime}}}$, and ${\mathbf{h_{g}^{\prime}}}$. (For $\mathbf{h_{b}}$, FC is applied row-wise.) Finally, a feature infusing step on $\mathbf{h_{a}}^{\prime}$ and $\mathbf{h_{b}}^{\prime}$ is performed:
@@ II-D Loss Function
III EXPERIMENTS
III-A Datasets and Labels
@@
GraphAU-Pain was trained on the UNBC-McMaster Shoulder Pain Expression Archive Database [19]. The dataset contains 48,398 colored frames from 25 participants with shoulder problems, showing facial expressions during pain-inducing actions. The faces were detected with OpenCV’s haarcascade frontalface default classifier and cropped to $172\times 172$. Each frame has 10 AU intensities ($0-5$) and PSPI scores ($0-16$), calculated as $PSPI=AU4+\max(AU6,AU7)+\max(AU9,AU10)+AU43$ [23]. PSPI pain intensity is categorized into ordinal levels: No Pain ($PSPI=0$), Mild Pain ($PSPI\in[1,4]$), and Obvious Pain ($PSPI\geq 5$). The categories are distributed in a skewed way, with each category respectively consisting approximately $82\%$, $15\%$ and $3\%$ of the UNBC dataset. This categorization is more clinically meaningful and interpretable compared to the method used in the GLA-CNN paper [26], as it mitigates the high sensitivity of PSPI scores to AU variations. Moreover, we replaced AU intensity with occurrence by capping the AU score at one.
III-B Training and Evaluation Details
@@
To prepare AU occurrence prediction before pain intensity estimation training, the supervised fine-tuning (SFT) process of GraphAU-Pain’s AU representation learning module was performed on the ANFL component with weights pretrained on DISFA for 20 epochs provided by Luo et al. [20]. Our relabeled UNBC+ dataset was used in the SFT process and undersampled to address data imbalance. The undersampling process involved randomly removing approximately $90\%$ of facial images with $PSPI=0$ and excluding facial images without active AUs. The full list of frames included in this subset is made available in the code repository. Through SFT for 17 epochs with a learning rate of $1e^{-5}$, a batch size of 16, and an Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a weight decay of $5e^{-4}$, the AU representation learning module achieved remarkable results in AU detection, compared to state-of-the-art results, as detailed in Table I. The actual training for pain intensity estimation used the same hyperparameters as the SFT process but was performed on the full original UNBC dataset.
TABLE I: F1-scores and accuracies achieved for each AU by the fine-tuned ANFL module.
AU1 AU2 AU4 AU6 AU9 AU12 AU25 AU26 Avg
@@
TABLE II: GraphAU-Pain’s performance statistics.
@@
Accuracy - 87.61
III-C Model Performance
Overall, GraphAU-Pain achieves a commendable average F1-score of $66.21\%$ and a high accuracy of $87.61\%$. Table II details the per-class results, showing strong performance for the No Pain category, which has an F1-score of $93.10\%$. However, performance declines for the Mild and Obvious categories, reflecting F1-scores of $51.19\%$ and $54.35\%$, respectively. This substantial difference is largely attributable to the dataset’s pronounced class imbalance. The log-scaled confusion matrix in Figure 2 further illustrates the distribution of predictions. While No Pain dominates the diagonal, indicating high accuracy there, some off-diagonal misclassifications occur between Mild and Obvious, and there is a noticeable bias toward predicting No Pain. Consequently, while the model is robust in detecting No Pain, additional strategies are needed to better distinguish between higher pain intensities.
@@ III-C Model Performance
Figure 2: Confusion matrix of GraphAU-Pain
The numbers are transformed in $\log_{10}$ scale for better visualization; the original number is shown below the scaled one.
III-D Comparison to SOTA
To the best of our knowledge, GLA-CNN [26] is the only other method that uses AUs for pain-intensity estimation on UNBC while focusing on cross-sectional facial image frames. No additional methods apply exactly the same categorization scheme, so we compare GraphAU-Pain with GLA-CNN and other models reported in [26]. To align labels, we reclassify pain intensities into four categories: No Pain ($PSPI=0$), Weak Pain ($PSPI=1$), Mild Pain ($PSPI=2$), and Strong Pain ($PSPI\geq 3$). By altering only the model’s final layer from 3 to 4 outputs and keeping other settings unchanged, GraphAU-Pain shows substantial gains in both accuracy and average F1-score, as shown in Table III and Figure 2(b). Note that GLA-CNN and the other compared models were trained and evaluated on an undersampled subset of UNBC [26] to deal with class imbalance, whereas GraphAU-Pain is trained on the full dataset. Therefore, their published performance might be higher than what would have been achieved on the full dataset.
TABLE III: Comparing classification performance of GraphAU-Pain with other models.
The last 3 models are retrained and tested on the same dataset as GLA-CNN, but only accuracy is reported [26].

### v2

@@ 1 Introduction
Recent advances in graph neural networks (GNNs), such as Multi-dimensional Edge Feature-based AU Relation Graph for AU (ME-GraphAU) [20] and Graph Relation Network (GRU) [21], have demonstrated promising AU prediction performance on the DISFA dataset [22] and BP4D dataset [23]. These methods use CNN backbones like ResNet [24] or VGG [25] to learn image features for each AU, then construct a relational graph based on the AU features. Each node in such graphs represents an AU, and each edge represents the relationship between a pair of AUs [26]. Through message propagation in GNN layers, the output AU features can capture individual AU information from neighbors as well as structural information [27]. Such graph-based AU features can then be aggregated to build a full-face representation informed by AUs and their relationships for downstream tasks [28], inspiring us to incorporate graph-based AU detection into pain intensity estimation.
Motivated by the above, we introduce GraphAU-Pain for accurate and interpretable pain intensity estimation and summarize our three contributions as follows:
  • GraphAU-Pain Model. We propose a novel graph-based framework that transforms AU detection into pain estimation by modeling AU relationships as a dynamic graph structure, enabling more expressive and interpretable pain assessment compared to traditional image-driven approaches.
  • Cross-Dataset Transfer with Relabeling. We introduce a novel transfer learning strategy that leverages DISFA-pretrained weights to address UNBC’s limited training data. By creating a hybrid UNBC+ dataset that combines original annotations with predicted labels for missing AUs, we enable effective knowledge transfer while preserving dataset integrity, significantly improving model performance in AU occurrence prediction.
@@
2 Related Work
2.1 Facial Action Unit Detection
FACS [13] categorizes facial expressions into 44 fundamental components known as AUs, each corresponding to specific muscle movements with intensity ratings from 0 to 5. For instance, AU1 represents the “Inner Brow Raiser” and AU2 the “Outer Brow Raiser.” These AUs provide a systematic method for analyzing facial expressions, including pain and other affective states. Early approaches to AU detection relied on traditional machine learning methods, with OpenFace [29] being a widely used open-source tool based on SVMs [30]. Deep learning methods like Deep Region and Multi-label Learning [31] established early benchmarks but failed to model AU interdependencies. Later attention-based methods [32, 33] improved performance significantly, though their pixel-wise approach faced limitations in capturing subtle facial changes and generalizing across diverse racial groups.
Graph Neural Networks (GNNs) have emerged as a powerful architecture for modeling complex dependencies between facial landmarks or AUs. A GNN layer passes information between nodes through edges, allowing nodes to learn their neighborhood information. This makes GNNs essential for learning relational representations, as demonstrated in fields like knowledge graph construction [34] and recommender systems [35]. In facial AU analysis, some studies introduced prior AU co-occurrence knowledge via Graph Convolutional Networks [36, 37], while others like Graph Relation Network (GRN) [21] explored knowledge-free approaches. GRN constructs a fully connected directed graph with image features as nodes and uses attention-based edge functions, achieving MAE of $0.7$ on BP4D and $0.2$ on DISFA. Unlike GRN, which fully connects all nodes, ME-GraphAU [20] introduces a novel approach by using a CNN backbone to extract features for each node that represents an AU and establishing connections between nodes based on feature similarity. This architecture effectively captures facial feature relationships by modeling AUs as nodes and their interactions as edges, achieving strong performance with F1 scores of $65.5\%$ on BP4D and $63.1\%$ on DISFA. Despite these advances, direct adaptation of AU detection models to pain estimation tasks remains challenging due to two key limitations. First, pain estimation datasets are typically much smaller than those used for AU detection. Second, these datasets exhibit a significant class imbalance in AU occurrences. Our GraphAU-Pain framework addresses both challenges through a novel transfer learning strategy, as detailed in Sec. 3.4.
2.2 Pain Prediction Based on Facial Expressions
Early approaches relied on feature extraction and classification techniques like KNN [16] and Random Forest [38]. These handcrafted feature approaches were limited by variations in head pose, lighting, and spontaneous expressions.
Recent deep learning approaches have evolved from traditional CNNs [39, 40] to more sophisticated architectures that incorporate AU features. While CNNs provide a foundation through pixel-level analysis, their lack of physiological knowledge limits both interpretability and generalizability. This limitation has driven the development of more advanced approaches, such as LSTM-based continuous pain monitoring [41]. However, this promising work faces practical constraints due to its reliance on a private dataset [42]. Similarly, while AU-based pain prediction has shown potential on BP4D [43], its binary classification approach may not capture the nuanced spectrum of natural pain experiences.
The state-of-the-art GLA-CNN [18] represents a significant advancement by combining CNNs with attention mechanisms to analyze facial pain and AU relationships. On the UNBC dataset, it achieves $36.2\%$ F1-score and $56.5\%$ accuracy. While these metrics show improvement, they remain insufficient for clinical applications. The model’s fine-grained category scheme for pain levels fails to account for PSPI’s sensitivity to AU intensity. This limitation becomes apparent when considering cases such as intense AU4 (brow lowering) can occur without pain [44], leading to confusion between its Weak Pain ($PSPI=0$), Weak Pain ($PSPI=1$), and Mild Pain ($PSPI=2$) categories. Furthermore, the black-box nature of the design limits clinical utility by obscuring how the estimated pain intensity is derived. These challenges highlight the need for more reliable and interpretable pain detection models, motivating our development of GraphAU-Pain. Our approach explicitly incorporates AUs through a graph-based modeling framework, providing enhanced interpretability through transparent AU-pain intensity relationships and improved accuracy via comprehensive modeling of AU interdependencies.
3 Material and Methods
3.1 Implementation Overview
@@
The GraphAU-Pain model is designed to estimate pain intensity based on facial image data, and its training involves two key steps. First, the full-face and AU representation learning modules are trained for AU occurrence prediction. These modules are adapted from the AU Relationship-aware Node Feature Learning (ANFL) component of ME-GraphAU [20]. Because directly training ANFL for AU occurrence prediction on the UNBC dataset only yielded a $20\%$ average F1-score, we developed a cross-dataset transfer learning strategy to improve performance. This strategy transfers AU prediction capabilities from the DISFA dataset model trained by Luo et al. [20] to the UNBC dataset. To address differences in AU labels between datasets, we created a relabeled UNBC+ dataset to align the AU annotations and used undersampling to address AU imbalance. Second, with the weights trained for AU prediction, GraphAU-Pain is then trained for pain intensity estimation with the full UNBC dataset. Facial images are first processed through the CNN backbone to extract pixel-wise full-face representations. These representations are then transformed into graph-based structures to learn AU-specific features. Finally, the full-face and AU-based features are combined to estimate pain intensity. The performance of the model is evaluated using two metrics: average F1-score and accuracy.
Figure 1: Illustration of the GraphAU-Pain model architecture. A global CNN backbone (orange) extracts a full-face representation, which is split into three feature branches by the AU representation learning module (green) and processed by the pain level classification module (blue). ANFL denotes the AU Relationship-aware Node Feature Learning component, FC the fully-connected layer, and MatMult matrix multiplication. GraphAU-Pain is trained by first initializing weights of the representation learning modules with the AU occurrence prediction task and then train the full model for the pain intensity estimation task.
@@
3.2 Model
@@
After acquiring the backbone feature $\mathbf{h_{b}}$, the AU representation module learns two distinct embeddings: $\mathbf{h_{a}}\in\mathbb{R}^{d_{\text{AU}}}$ ($d_{\text{AU}}=512$) encoding individual AU occurrences, and $\mathbf{h_{g}}\in\mathbb{R}^{d_{\text{AU}}}$ encoding the complete AU relational graph structure. The module first transforms $\mathbf{h_{b}}$ through $n_{\text{AU}}$ fully connected layers to generate initial AU representations $\mathbf{H_{a}}\in\mathbb{R}^{n_{\text{AU}}\times d_{\text{AU}}}$, where each row represents one AU. These representations serve as node features in a graph where each node connects to its $K=3$ most similar nodes based on dot-product similarity. The graph structure is then processed through a graph convolutional layer:
$\mathbf{H_{a}^{\prime}}\;=\;\text{ReLU}\Bigl{(}\mathbf{H_{a}}+\text{BN}\bigl{(}\mathbf{A}^{\top}\text{FC}_{1}(\mathbf{H_{a}})+\text{FC}_{2}(\mathbf{H_{a}})\bigr{)}\Bigr{)}\;\in\;\mathbb{R}^{n_{\text{AU}}\times d_{\text{AU}}},$ (1)
where BN denotes batch normalization, FC represents fully connected layers, and $\mathbf{A}$ is the normalized adjacency matrix. The graph representation $\mathbf{h_{g}}\in\mathbb{R}^{d_{\text{AU}}}$ is obtained through global sum pooling of the node embeddings. Although we could also add an edge update module here [20], we omitted it to avoid overfitting on the UNBC dataset. For each AU, its occurrence probability $p_{i}$ is computed as the cosine similarity between its representation $h_{a,i}$ and a learnable vector $s_{i}$:
$p_{i}=\frac{\text{ReLU}(h_{a,i})^{\top}\text{ReLU}(s_{i})}{\|\text{ReLU}(h_{a,i})\|_{2}\|\text{ReLU}(s_{i})\|_{2}},$ (2)
[…]
@@ 3.3 Loss Function
3.4 Transfer Learning
Our preliminary experimental results showed that directly training GraphAU-Pain on UNBC yielded unsatisfactory results, with 30–40% F1-score and 60–80% accuracy after trying several settings. To tackle this problem and improve AU prediction performance through transfer learning, in this work we initialized the weights of the full-face and AU representation modules with the weights pretrained on DISFA provided by Luo et al. [20]. We chose the DISFA dataset [22] for pretraining because it is three times larger than UNBC and provides high-quality AU annotations that align well with UNBC’s AU annotation scheme—sharing six out of eight AUs with UNBC and three of them are used in PSPI calculation—making this dataset suitable for transfer learning. However, since DISFA includes two additional AUs (AU1 and AU2) not present in UNBC’s original annotations, for fine-tuning the pretrained weights, we need to label these additional AUs for UNBC to ensure complete AU coverage. To do this, we pass UNBC’s facial images through the pretrained representation learning modules to predict all eight AU labels. We then create a hybrid dataset (UNBC+) by keeping UNBC’s original annotations for the six overlapping AUs while using the predicted values for AU1 and AU2. This relabeling process ensures that our model can learn from a complete set of AU activations while maintaining the reliability of UNBC’s original annotations where available.
4 Experiments
4.1 Datasets and Labels
@@
GraphAU-Pain was trained on the UNBC-McMaster Shoulder Pain Expression Archive Database [15]. The dataset contains 48,398 colored frames from 25 participants with shoulder problems, showing facial expressions during pain-inducing actions. The faces were detected with OpenCV’s haarcascade frontalface default classifier and cropped to $172\times 172$. Each frame has 10 AU intensities ($0-5$) and PSPI scores ($0-16$), calculated as $PSPI=AU4+\max(AU6,AU7)+\max(AU9,AU10)+AU43$ [14]. PSPI pain intensity is categorized into ordinal levels: No Pain ($PSPI=0$), Mild Pain ($PSPI\in[1,4]$), and Obvious Pain ($PSPI\geq 5$). The categories are distributed in a skewed way, with each category respectively consisting approximately $82\%$, $15\%$, and $3\%$ of the UNBC dataset. This categorization is more clinically meaningful and interpretable compared to the method used by Wu et al. [18], as it mitigates the high sensitivity of PSPI scores to AU variations. Moreover, we replaced AU intensity with occurrence by capping the AU score at one.
4.2 Training and Evaluation Details
@@
To prepare AU occurrence prediction before pain intensity estimation training, the supervised fine-tuning (SFT) process of GraphAU-Pain’s AU representation learning module was performed on the ANFL component with weights pretrained on DISFA for 20 epochs provided by Luo et al. [20]. Our relabeled UNBC+ dataset was used in the SFT process and undersampled to address data imbalance. The undersampling process involved randomly removing approximately $90\%$ of facial images with $PSPI=0$ and excluding facial images without active AUs. The full list of frames included in this subset is made available in the code repository. We use all $n_{\text{AU}}=8$ AU labels in the UNBC+ dataset as listed in Table 1. We trained the module through SFT for 17 epochs with a learning rate of $1e^{-5}$, a batch size of 16, and an Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a weight decay of $5e^{-4}$. The AU representation learning module achieved remarkable results in AU detection, compared to state-of-the-art results, as detailed in Table 1. The training for pain intensity estimation used the same hyperparameters as the SFT process but was performed on the full original UNBC dataset.
Table 1: F1-scores and accuracies achieved for each AU by the fine-tuned representation learning module. The average F1-score and accuracy are calculated by averaging the F1-scores and accuracies of all AUs. While accuracy remains high across all AUs, F1-scores are notably lower for AU9 and AU26, likely due to class imbalance.
Metric AU1 AU2 AU4 AU6 AU9 AU12 AU25 AU26 Avg
@@
Table 2: GraphAU-Pain’s performance statistics. Our model achieves strong performance for No Pain detection but lower and similar performance for Mild and Obvious pain categories, likely due to data imbalance. The overall F1/Recall/Precision/Accuracy are calculated by averaging the F1/Recall/Precision/Accuracy of all categories.
@@
Accuracy - - - 87.61
4.3 Model Performance
Overall, GraphAU-Pain achieves a commendable average F1-score of $66.21\%$ and a high accuracy of $87.61\%$. Table 2 details the per-class results, showing strong performance for the No Pain category, which has an F1-score of $93.10\%$. However, performance declines for the Mild and Obvious categories, reflecting F1-scores of $51.19\%$ and $54.35\%$, respectively. This performance gap between No Pain and the Mild/Obvious categories is largely attributable to the dataset’s pronounced class imbalance. The log-scaled confusion matrix in Figure 2 further illustrates the distribution of predictions. While No Pain dominates the diagonal, indicating high accuracy there, some off-diagonal misclassifications occur between Mild and Obvious, and there is a noticeable bias toward predicting No Pain. Consequently, while the model is robust in detecting No Pain, additional strategies are needed to better distinguish between higher pain intensities.
@@ 4.3 Model Performance
Figure 2: Confusion matrix of GraphAU-Pain. The numbers are transformed in $\log_{10}$ scale for better visualization; the original number is shown below the scaled one. The 3-category classification (left) is introduced in this work to mitigate the excessive sensitivity of the 4-category system (right) proposed by Wu et al. [18]. The 4-category system’s additional “Weak Pain” category (PSPI=1) proves too sensitive to AU intensity variations and lacks clinical significance when separated from our “Mild Pain” category. Therefore, the lower performance in recognizing “Weak Pain” reflects the metric’s sensitivity to minor AU variations rather than model limitations.
4.4 Comparison to SOTA
To the best of our knowledge, GLA-CNN [18] is the only other method that uses AUs for pain-intensity estimation on UNBC while focusing on cross-sectional facial image frames. No additional methods apply exactly the same categorization scheme, so we compare GraphAU-Pain with GLA-CNN and other models reported in [18]. To align labels, we reclassify pain intensities into four categories: No Pain ($PSPI=0$), Weak Pain ($PSPI=1$), Mild Pain ($PSPI=2$), and Strong Pain ($PSPI\geq 3$). By altering only the model’s final layer from three to four outputs and keeping other settings unchanged, GraphAU-Pain shows substantial gains in both accuracy and average F1-score, as shown in Table 3 and Figure 2(b). Note that GLA-CNN and the other compared models were trained and evaluated on an undersampled subset of UNBC [18] to deal with class imbalance, whereas GraphAU-Pain is trained on the full dataset. Therefore, their published performance might be higher than what would have been achieved on the full dataset.
Table 3: Comparison of GraphAU-Pain with other models on the UNBC dataset. We compare our results with those reported by Wu et al. [18], and we retrained our model with the 4-category system for fair comparison.